OpenStudy (anonymous):

A pole that is 10 feet tall casts a shadow 15 feet long. At the same time, another pole casts a shadow having a length of 12 feet. What is the height, in feet, of the second pole? Answer 18 12.5 9 8

OpenStudy (campbell_st):

use ratios of similar triangles x/10 = 12/15 x = (10 x 12)/15

OpenStudy (anonymous):

so 8 right

OpenStudy (campbell_st):

looks like it